# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ III.1 Fourier transform of maps
If instead of the complex valued function $f:G\rightarrow\mathbb{C}$, we take a map $\Phi:G\rightarrow\mathcal{A}$, where $\mathcal{A}$ is an arbitrary algebra, then it is natural to define the Fourier transform of $\Phi$ w.r.t a representation $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ as follows:
@@
where $\operatorname{tr}_{d_{\rho}}$ represents partial trace over $M_{d_{\rho}}(\mathbb{C})$ and $\mathrm{Irr}(G)$ represents the complete set of inequivalent irreducible representations of $G$. The Plancherel formula is given as follows( See Appendix (A) for a proof):
@@
Using the definition of convolution used in Theorem (A), one can prove that the Fourier transform of the convolution of two maps on a finite group is the product of their respective Fourier transform.
III.1.1 Fourier transform of maps from a contracted algebra to an arbitrary algebra
Let $S$ be a finite semigroup and $z\in S$ be the zero element. We define the Fourier transform of a linear map $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ w.r.t a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$ as
@@ III.1.1 Fourier transform of maps from a contracted algebra to an arbitrary algebra
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of a semigroup $S$ and $\mathcal{A}$ be an arbitrary algebra. Then for maps $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ and $\Phi^{\prime}:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$, and a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$ we have
@@
Now one can ask if there is a one-to-one correspondence between $\Phi$ and $\widehat{\Phi}(\rho)$. In general there is no such correspondence, however if we consider a complete set of irreducible representations $\rho$ of $\mathbb{C}_{0}[S]$, then Wedderburn-Artin theorem helps us to built a one-to-one correspondence. The inverse semigroup algebra $\mathbb{C}[S]$ is semisimple [14]. Now it can be proved that $\mathbb{C}_{0}[S]$ is semisimple if and only if $\mathbb{C}[S]$ is semisimple [3]. Wedderburn isomorphism states that $\mathbb{C}_{0}[S]$ is isomorphic to $\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})$, with $\mathrm{Irr}(\mathbb{C}_{0}[S])$ being the complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$. The semigroup algebra $\mathbb{C}_{0}[S]$ is isomorphic to the convolution algebra $L(\mathbb{C}_{0}[S],\mathbb{C})$ implying that $L(\mathbb{C}_{0}[S],\mathbb{C})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})$. In particular, the isomorphism is given by $\Phi\mapsto\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\left(\sum_{s\in S\setminus\{0\}}\Phi(s)\rho(s)\right)$. This simple observation generalizes to $L(\mathbb{C}_{0}[S],\mathcal{A})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})\otimes\mathcal{A}$ with $\Phi\mapsto\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\left(\sum_{s\in S\setminus\{0\}}\rho(s)\otimes\Phi(s)\right)$. Equivalently, we can say that there is one-to-one correspondence between $\Phi$ and $\bigoplus_{\rho\in\mathcal{X}}\widehat{\Phi}(\rho)$.
III.2 Fourier Inversion for finite inverse semigroup
In this subsection, we will derive the Fourier inversion formula for finite inverse semigroup. The main idea behind the proof is the following: Steinberg’s isomorphism provides a way to construct a complete set of inequivalent irreducible representations of the contracted algebra of a finite inverse semigroup in terms of the complete set or inequivalent irreducible representations of the maximal subgroups of the $D$-classes of the inverse semigroup (see Theorem (J)). Then the Fourier inversion formula for the maximal subgroups helps to provide the Fourier inversion for the inverse semigroup. The Fourier inversion formula for complex valued functions on the semigroup algebra of a finite inverse semigroup is derived in [15]. Here, following a similar technique, we derive the Fourier inversion formula for matrix-valued functions on the contracted algebra of a finite inverse semigroup.
@@ III.2 Fourier Inversion for finite inverse semigroup
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of an inverse semigroup $S$. Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ denotes a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$. Let $\{D_{k}\}_{k=1}^{n}$ represents the set of $\mathcal{D}$-classes of $S$, $r_{k}$ denotes the number of idempotent elements in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup at an arbitrarily chosen idempotent element $e_{k}\in D_{k}$. Then the inversion formula for the Fourier transform $\widehat{\Phi}(\rho)=\sum_{s\in S,s\neq 0}\rho(s)\otimes\Phi(s)=\sum_{s\in S,s\neq 0}\rho(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor)$ of a linear map $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is given by
@@
where $d_{\sigma}$ is the dimension of the representation $\sigma$ and the linear map $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is related to the map $\Phi$ by $\Phi(s)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\mu(s,t)\widetilde{\Phi}(\lfloor t\rfloor)$ and $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$.
Proof: We denote the complete set of irreducible representations of $\mathbb{C}_{0}[S]$ constructed in the way mentioned in Theorem (J) by $\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$. Specifically, the representation $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$ is given by the following equation:
@@
Notice that $\sum_{s\in S,s\neq 0}s\otimes\Phi(s)\in\mathbb{C}_{0}[S]\otimes M_{n}(\mathbb{C})$ can also be expressed in terms of the groupoid basis as $\sum_{s\in S,s\neq 0}\lfloor s\rfloor\otimes\widetilde{\Phi}(\lfloor s\rfloor)$. The linear maps $\widetilde{\Phi}(\lfloor s\rfloor)$ and $\Phi(s)$ are related to each other by the following relations
@@
As the Fourier transform of $\Phi$ is $\widehat{\Phi}(\rho)=(\rho\otimes\mathrm{id})\sum_{s\in S,s\neq 0}s\otimes\Phi(s)$, we have
@@
Now for $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$,
@@
Notice that the Fourier transform $\widehat{\Phi}(\bar{\rho}_{k})$ is a matrix indexed by the idempotents of $S$. For idempotents $a,b\in D_{k}$, the $(a,b)$-th element of $\widehat{\Phi}(\bar{\rho}_{k})$ is given by
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}s\in D_{k}\\ \mathrm{ran}(s)=a\\ \mathrm{dom}(s)=b\end{subarray}}\rho_{k}(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)})\otimes\widetilde{\Phi}(\lfloor s\rfloor).$ (43)
We notice from Proposition (H) that $u:=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}$ is an element of $G_{e_{k}}$. We can invert the relation $u=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}$ as follows: $u=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=p_{\mathrm{ran}(s)}p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}p^{-1}_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=\mathrm{ran}(s)\,s\,\mathrm{dom}(s)=ss^{-1}ss^{-1}s=s$. As $\mathrm{ran}(s)=a$ and $\mathrm{dom}(s)=b$, we have, $s=p_{a}\,u\,p^{-1}_{b}$. With this observation, we can write the summation in Eq. (43) as sum over those elements $u\in G_{e_{k}}$ such that $\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a$ and $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$:
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}u\in G_{e_{k}}\\ :\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a,\\ \mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b\end{subarray}}\rho_{k}(u)\otimes\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor).$ (44)
Using Proposition (I), we can rewrite Eq. (44) as follows:
@@
We can immediately identify $\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ in Eq. (45) as the Fourier transformation of the map $\widetilde{\Phi}(\lfloor p_{a}\,(\cdot)\,p^{-1}_{b}\rfloor)$ on the maximal subgroup $G_{e_{k}}$. Hence, we use Eq. (33) for the Fourier inversion and arrive at the following:
$\widetilde{\Phi}(\lfloor s\rfloor)=\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor)$ $=\frac{1}{|G_{e_{k}}|}\sum_{\rho_{k}\in\mathrm{Irr(G)}}d_{\rho_{k}}\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]},$ (46)
where $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. From Eq. (42), we have
@@
where $i$ and $j$ are idempotents in $D_{k}$. A quick calculation shows that
$\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]}$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big{[}(E_{b,a}\otimes\rho_{k}(u^{-1})\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big{]}.$ (48)
As $s\in D_{k}$, from Eq. (38), we have
@@
$\operatorname{tr}_{d_{\rho_{k}}}\big{[}(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\big{]}$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big{[}(\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big{]}.$ (51)
Notice that $\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}}=\operatorname{tr}_{d_{\overline{\rho}_{k}}}$. Using Eq. (51) in Eq. (46), we have
@@
Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ be an arbitrary complete set of inequivalent irreducible representation of $\mathbb{C}_{0}[S]$. Let $\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])$. Then $\sigma$ must be equivalent to an element $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$, i.e there exists an invertible transformation $H$ such that $\overline{\rho}_{k}(s)=H^{-1}\sigma(s)H$ for all $s\in S$. From Eq. (41), we have
@@
We notice that by construction, the dimension of the induced representation $d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$. As $\overline{\rho}_{k}$ is equivalent to the representation $\sigma$, their dimension must be the same, i.e., $d_{\sigma}=d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$ implying $d_{\rho_{k}}=\frac{d_{\sigma}}{r_{k}}$. With this observation, we have the following inversion formula
@@
Now we focus on the finite dimensional matrix algebra $M_{m}(\mathbb{C})$. It can be verified that the set $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$ of matrix unit along with the zero matrix form an inverse semigroup and the matrix algebra $M_{m}(\mathbb{C})$ is the contracted semigroup algebra, i.e., $\mathbb{C}_{0}[S_{M}]=M_{m}(\mathbb{C})$. The idempotent elements of $S_{M}$ are $e_{ii}$ for all $i\in\{1,2,...,m\}$ (and of course 0). The inverse of an element $e_{ij}$ is the $e_{ji}$ i.e., $e_{ij}^{-1}=e_{ji}$. It can also be verified that the maximal subgroup at an idempotent $e_{ii}$ is $G_{e_{ii}}=\{e_{ii}\}$.
@@
The $D$-classes of the inverse semigroup $S_{M}$ are $D_{0}:=\{0\}$ and $D_{1}:=\{e_{ij}\}_{i,j=1}^{m}$.
Proof: Let $e_{ij}\mathcal{D}~0$. Then there exist $e_{pq}$ such that $e_{qp}e_{pq}=0$ and $e_{pq}e_{qp}=e_{ij}e_{ji}$. However For a matrix unit $e_{pq}$, we always have $e_{qp}e_{pq}\neq 0$. Hence the $D$-class containing the 0``0" matrix has the element 0``0" only. Now, we have the element $e_{ki}$ which satisfies $e_{ki}^{-1}e_{ki}=e_{ij}e_{ij}^{-1}$ and $e_{ki}e_{ki}^{-1}=e_{kl}e_{kl}^{-1}$. Hence we have $e_{ij}\mathcal{D}e_{kl}$ for arbitrary $i,j,k,l\in\{1,2,...,m\}$. This completes the proof. $\Box$

### v2

@@ 3.1 Fourier Transform of Maps
If, instead of the complex valued function $f:G\rightarrow\mathbb{C}$, we consider a map $\Phi:G\rightarrow\mathcal{A}$, where $\mathcal{A}$ is an arbitrary algebra, it is natural to define the Fourier transform of $\Phi$ with respect to a representation $\rho:G\rightarrow M_{d_{\rho}}(\mathbb{C})$ as follows:
@@
where $\operatorname{tr}_{d_{\rho}}$ denotes partial trace over $M_{d_{\rho}}(\mathbb{C})$, and $\mathrm{Irr}(G)$ denotes a complete set of inequivalent irreducible representations of $G$. The Plancherel formula is given by (See Appendix A for a proof):
@@
Using the definition of convolution used in Theorem A, one can prove that the Fourier transform of the convolution of two maps on a finite group is the product of their respective Fourier transform.
3.1.1 Fourier Transform for Linear Maps on Contracted Semigroup Algebras
Let $S$ be a finite semigroup and $z\in S$ be the zero element. The Fourier transform of a linear map $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ with respect to a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$ is defined as
@@ 3.1.1 Fourier Transform for Linear Maps on Contracted Semigroup Algebras
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of a semigroup $S$ and $\mathcal{A}$ be an arbitrary algebra. Then for maps $\Phi:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$ and $\Phi^{\prime}:\mathbb{C}_{0}[S]\rightarrow\mathcal{A}$, and a representation $\rho:\mathbb{C}_{0}[S]\rightarrow M_{d_{\rho}}(\mathbb{C})$, we have
@@
Now we may ask whether there exists a one-to-one correspondence between $\Phi$ and $\widehat{\Phi}(\rho)$. In general, such a correspondence does not exist. However, if $S$ is a finite inverse semigroup, and one considers a complete set of irreducible representations $\rho$ of $\mathbb{C}_{0}[S]$, then Wedderburn-Artin theorem yields such a correspondence.
The inverse semigroup algebra $\mathbb{C}[S]$ is semisimple [37], and $\mathbb{C}_{0}[S]$ is semisimple if and only if $\mathbb{C}[S]$ is semisimple [26]. Consequently there is an algebra isomorphism
$\mathbb{C}_{0}[S]\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C}),$
where $\mathrm{Irr}(\mathbb{C}_{0}[S])$ denotes a complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$. Identifying $\mathbb{C}_{0}[S]$ with the convolution algebra $L(\mathbb{C}_{0}[S],\mathbb{C})$, one obtains
$L(\mathbb{C}_{0}[S],\mathbb{C})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C}),$
with the explicit correspondence
$\Phi\mapsto\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\sum_{s\in S\setminus\{0\}}\Phi(s)\rho(s).$
This simple observation extends naturally to $\mathcal{A}$-valued functions, yielding
$L(\mathbb{C}_{0}[S],\mathcal{A})\cong\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}M_{d_{\rho}}(\mathbb{C})\otimes\mathcal{A}$ (44)
with the correspondence
$\Phi\mapsto\!\!\!\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\sum_{s\in S\setminus\{0\}}\rho(s)\otimes\Phi(s).$ (45)
Equivalently, one can say that there is one-to-one correspondence between $\Phi$ and $\bigoplus_{\rho\in\mathrm{Irr}(\mathbb{C}_{0}[S])}\widehat{\Phi}(\rho)$.
3.2 Fourier Inversion for Finite Inverse Semigroups
In this section, we derive a Fourier inversion formula for finite inverse semigroup. The proof relies on the following idea: Steinberg’s isomorphism provides a way to construct a complete set of inequivalent irreducible representations of the contracted algebra of a finite inverse semigroup in terms of complete sets of inequivalent irreducible representations of the maximal subgroups associated with the $\mathcal{D}$-classes of the inverse semigroup (see Theorem J). The Fourier inversion formula for these maximal subgroups then helps to provide the Fourier inversion for the inverse semigroup. While the Fourier inversion formula for complex valued functions on the semigroup algebra of a finite inverse semigroup was established in [38], we extend this approach to derive a Fourier inversion formula for matrix-valued functions on the contracted algebra of a finite inverse semigroup.
@@ 3.2 Fourier Inversion for Finite Inverse Semigroups
Let $\mathbb{C}_{0}[S]$ be the contracted semigroup algebra of a finite inverse semigroup $S$. Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ denote a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$. Let $\{D_{k}\}_{k=1}^{n}$ be the set of $\mathcal{D}$-classes of $S$, let $r_{k}$ be the number of idempotents in $D_{k}$, and let $G_{e_{k}}$ be the maximal subgroup at an arbitrarily chosen idempotent element $e_{k}\in D_{k}$. Then the inversion formula for the Fourier transform
$\widehat{\Phi}(\rho)=\sum_{s\in S,s\neq 0}\rho(s)\otimes\Phi(s)=\sum_{s\in S,s\neq 0}\rho(\lfloor s\rfloor)\otimes\widetilde{\Phi}(\lfloor s\rfloor)$
of a linear map $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ is given by
@@
where $d_{\sigma}$ denote the dimension of the representation $\sigma$. The linear maps $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ and $\Phi$ are related by $\Phi(s)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\mu(s,t)\widetilde{\Phi}(\lfloor t\rfloor)$, and $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$, where $\mu$ denotes the Möbius function on the natural partial order of $S$.
Proof: Let $\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$ denote the complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ constructed as in Theorem J. More precisely, a representation $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$ is given by
@@
Notice that the element $\sum_{s\in S,s\neq 0}s\otimes\Phi(s)\in\mathbb{C}_{0}[S]\otimes M_{n}(\mathbb{C})$ can also be expressed in terms of the groupoid basis as $\sum_{s\in S,s\neq 0}\lfloor s\rfloor\otimes\widetilde{\Phi}(\lfloor s\rfloor)$. The linear maps $\widetilde{\Phi}$ and $\Phi$ are related to each other by the following relations:
@@
Since the Fourier transform of $\Phi$ can be written as $\widehat{\Phi}(\rho)=(\rho\otimes\mathrm{id})\sum_{s\in S,s\neq 0}s\otimes\Phi(s)$, it follows that
@@
Now, for $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$,
@@
Observe that the Fourier transform $\widehat{\Phi}(\bar{\rho}_{k})$ is a matrix indexed by the idempotents of $D_{k}$. For idempotents $a,b\in D_{k}$, the $(a,b)$-th entry of $\widehat{\Phi}(\bar{\rho}_{k})$ is given by
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\sum_{\begin{subarray}{c}s\in D_{k}\\ \mathrm{ran}(s)=a\\ \mathrm{dom}(s)=b\end{subarray}}\rho_{k}\left(p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}\right)\otimes\widetilde{\Phi}(\lfloor s\rfloor).$ (52)
By Proposition G, the element $u:=p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}$ is an element of the maximal subgroup $G_{e_{k}}$ at an idempotent $e_{k}\in D_{k}$. We can invert the relation $u=p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}$ as follows: $u=p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=p_{\mathrm{ran}(s)}p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}p^{-1}_{\mathrm{dom}(s)}\iff p_{\mathrm{ran}(s)}up^{-1}_{\mathrm{dom}(s)}=\mathrm{ran}(s)\,s\,\mathrm{dom}(s)=ss^{-1}ss^{-1}s=s$. Since $\mathrm{ran}(s)=a$ and $\mathrm{dom}(s)=b$, we have, $s=p_{a}\,u\,p^{-1}_{b}$. With this observation, we can write the summation in Eq. (52) as sum over those elements $u\in G_{e_{k}}$ such that $\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a$ and $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$:
$\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ $=\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}u\in G_{e_{k}}\\ \mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a,\\ \mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\!\rho_{k}(u)\otimes\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor).$ (53)
By Proposition H, we can rewrite Eq. (53) as follows:
@@
We can immediately identify $\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}$ in Eq. (54) as the Fourier transform of the map $\widetilde{\Phi}(\lfloor p_{a}\,(\cdot)\,p^{-1}_{b}\rfloor)$ on the maximal subgroup $G_{e_{k}}$. Hence, we use Eq. (40) for the Fourier inversion and arrive at the following:
$\widetilde{\Phi}(\lfloor s\rfloor)=\widetilde{\Phi}(\lfloor p_{a}\,u\,p^{-1}_{b}\rfloor)$ $=\frac{1}{|G_{e_{k}}|}\sum_{\rho_{k}\in\mathrm{Irr(G)}}d_{\rho_{k}}\operatorname{tr}_{d_{\rho_{k}}}\left[(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\right],$ (55)
where $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. It follows from Eq. (51) that
@@
where $i,j\in D_{k}$ are idempotents. A quick calculation shows
$\operatorname{tr}_{d_{\rho_{k}}}\left[(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\right]$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big[(E_{b,a}\otimes\rho_{k}(u^{-1})\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big].$ (57)
Since $s\in D_{k}$, from Eq. (47), we have
@@
$\operatorname{tr}_{d_{\rho_{k}}}\left[(\rho_{k}(u^{-1})\otimes\mathbb{I})\left(\widehat{\Phi}(\bar{\rho}_{k})\right)_{a,b}\right]$ $=(\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}})\big[(\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\bar{\rho}_{k})\big].$ (60)
Notice that $\operatorname{tr}_{r_{k}}\otimes\operatorname{tr}_{d_{\rho_{k}}}=\operatorname{tr}_{d_{\overline{\rho}_{k}}}$. Using Eq. (60) in Eq. (55), we have the following.
@@
Let $\mathrm{Irr}(\mathbb{C}_{0}[S])$ be an arbitrary complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$. Let $\sigma\in\mathrm{Irr}(\mathbb{C}_{0}[S])$. Then $\sigma$ must be equivalent to an element $\overline{\rho}_{k}\in\overline{\mathrm{Irr}}(\mathbb{C}_{0}[S])$, i.e., there exists an invertible transformation $H$ such that $\overline{\rho}_{k}(s)=H^{-1}\sigma(s)H$ for all $s\in S$. From Eq. (50), we have
@@
Notice that by construction, the dimension of the induced representation $d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$. Since $\overline{\rho}_{k}$ is equivalent to the representation $\sigma$, their dimensions must be the same, i.e., $d_{\sigma}=d_{\overline{\rho}_{k}}=r_{k}d_{\rho_{k}}$ and hence $d_{\rho_{k}}=\frac{d_{\sigma}}{r_{k}}$. With this observation, we obtain the following inversion formula
@@
Now we consider the finite dimensional matrix algebra $M_{m}(\mathbb{C})$. It can be verified that the set $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$ consisting of matrix units together with the zero matrix, forms an inverse semigroup. Moreover, the matrix algebra $M_{m}(\mathbb{C})$ is naturally identified with the contracted semigroup algebra of $S_{M}$, i.e., $\mathbb{C}_{0}[S_{M}]=M_{m}(\mathbb{C})$ [25]. The idempotent elements of $S_{M}$ are $e_{ii}$ for all $i\in\{1,2,...,m\}$ and 0. The inverse of an element $e_{ij}$ is the $e_{ji}$ i.e., $e_{ij}^{-1}=e_{ji}$. It can also be verified that the maximal subgroup at an idempotent $e_{ii}$ is $G_{e_{ii}}=\{e_{ii}\}$.
@@
The $\mathcal{D}$-classes of the inverse semigroup $S_{M}$ are $D_{0}:=\{0\}$ and $D_{1}:=\{e_{ij}\mid 1\leq i,j\leq m\}$.
Proof: Suppose that $e_{ij}\mathcal{D}~0$. Then there exist $e_{pq}$ such that $e_{qp}e_{pq}=0$ and $e_{pq}e_{qp}=e_{ij}e_{ji}$. However, for any matrix unit $e_{pq}$, one always has $e_{qp}e_{pq}\neq 0$, which yields a contradiction. Hence, the $D$-class containing the zero matrix consists only of 0.
Next, for arbitrary indices $i,j,k,l\in\{1,2,...,m\}$, consider the element $e_{ki}$. We have $e_{ki}^{-1}e_{ki}=e_{ij}e_{ij}^{-1}$ and $e_{ki}e_{ki}^{-1}=e_{kl}e_{kl}^{-1}$. Therefore, $e_{ij}\mathcal{D}e_{kl}$. This shows that all non-zero elements $e_{ij}$ lie in the same $\mathcal{D}$-class. $\Box$